# Two-dimensional delta Bose gas in a weighted space

### v1

@@
Abstract.
We extend the construction of the semigroup of the two-dimensional delta-Bose gas in [20] (based on [22, 16]) to a weighted $L^{2}$ space that allows exponentially growing functions. We further show that the semigroup of the mollified delta-Bose gas converges strongly to that of the delta-Bose gas.
1. Introduction
In this paper, we study the two-dimensional, $n$-particle delta-Bose gas. Fix a nonnegative $\Phi\in C^{\infty}(\mathbb{R}^{2})$ with a compact support such that $\int_{\mathbb{R}^{2}}\mathrm{d}x_{1}\,\Phi(x_{1})=1$, and, for $\varepsilon\to 0$, let $\Phi^{\varepsilon}(x_{1}):=\varepsilon^{-2}\Phi(\varepsilon^{-1}x_{1})$ be the corresponding approximation of the delta function. Let $[\![n]\!]:=\{1,\ldots,n\}$, write $\alpha=ij=\{i<j\}$ for an unordered pair of integers in $[\![n]\!]$, view $ij$ as a set, and let $\mathrm{Pair}[\![n]\!]:=\{ij|ij\subset[\![n]\!]\}$. Consider $\mathbb{R}^{2[\![n]\!]}:=(\mathbb{R}^{2})^{[\![n]\!]}=\{x=(x_{1},\ldots,x_{n})|x_{i}\in\mathbb{R}^{2}\}$, and, for $\alpha=ij$, let $\Phi^{\varepsilon}_{\alpha}(x):=\Phi^{\varepsilon}(x_{i}-x_{j})$. The Hamiltonian $\mathcal{H}^{\varepsilon}$ of the mollified delta-Bose gas acts on functions on $\mathbb{R}^{2[\![n]\!]}$ and is given by
$-\mathcal{H}^{\varepsilon}:=\frac{1}{2}\sum_{i=1}^{n}\Delta_{i}+\beta_{\varepsilon}\sum_{\alpha\in\mathrm{Pair}[\![n]\!]}\Phi^{\varepsilon}_{\alpha}\ ,$ (1.1)
@@ 1. Introduction
$\beta_{\varepsilon}$ $:=\frac{2\pi}{|\log\varepsilon|}+\frac{\pi}{|\log\varepsilon|^{2}}\Big{(}\theta-2\log 2+2\gamma+2\int_{\mathbb{R}^{4}}\mathrm{d}x_{1}\mathrm{d}x_{1}^{\prime}\Phi(x_{1})\log|x_{1}-x_{1}^{\prime}|\Phi(x_{1}^{\prime})\Big{)}\ ,$ (1.2)
@@
$\gamma=0.577\ldots$ denotes the Euler–Mascheroni constant, and $\theta$ is a fixed parameter that can take any real value. In (1.2), the constants after $\theta$ are just conventional (and can be absorbed into $\theta$).
@@
The delta-Bose gas is an instance of quantum many-body systems and is highly relevant in the study of the Stochastic Heat Equation (SHE). In the quantum context, (1.1) can be used to approximate the Hamiltonian of $n$ Bosons with pairwise, attractive (true) delta potentials. Quantum systems with delta potentials exhibit intriguing physical properties, post significant mathematical challenges, and enjoy a long history of study. We point to [2] for a review and mention the works [15, 22, 16, 19, 18] on the two-dimensional delta-Bose gas. As it turns out, in two dimensions, in order to see a non-degenerate limit as $\varepsilon\to 0$, one needs to attenuate the coupling constant $\beta_{\varepsilon}$ logarithmically around the value $2\pi$ as in (1.2). This was observed in the $n=2$ case in [2, Chapter I.5] and [3]. In the SHE context, the semigroup $\mathcal{Q}^{\varepsilon}(t):=e^{-t\mathcal{H}^{\varepsilon}}$ describes the moments of the noise-mollified SHE. Based on [2, 1], the work [3] established the convergence (as $\varepsilon\to 0$) of the second moment of the noise-mollified SHE to an explicit limit. Convergence of higher moments are significantly more challenging, and was later obtained in [5] for the third moment based on [4], and in [20] for of all moments based on [22, 16]; see [8, 9, 10] for probabilistic studies of the moments. A central question in this context concerns the $\varepsilon\to 0$ limit of the noise-mollified SHE, and the limit is called the Stochastic Heat Flow (SHF). The construction of the SHF had been a challenging open problem, which was accomplished in [6] by showing that the limit of the directed polymers is unique and universal. A different, axiomatic approach based the moments was recently carried out in [23].
@@
In this paper, we generalize the results of [20] from the $L^{2}$ space to a weighted $L^{2}$ space. Based on [22, 16], the work [20] gave an explicit formula of a strongly continuous semigroup $\mathcal{Q}(t)$ on $L^{2}(\mathbb{R}^{2[\![n]\!]})$ and showed that $\mathcal{Q}^{\varepsilon}(t):=e^{-t\mathcal{H}^{\varepsilon}}$ converges strongly to $\mathcal{Q}(t)$ on $L^{2}(\mathbb{R}^{2[\![n]\!]})$. These results translate into the convergence of the moments of the noise-mollified SHE to an explicit limit, for $L^{2}$ initial conditions. The class of $L^{2}$ functions is natural, because the generator of $\mathcal{Q}(t)$ is self-adjoint, but the class leaves out some commonly consider initial conditions, most notably the constant/flat ones. Further, when using delta-Bose gas to study the noise-mollified SHE and the SHF, one often needs to consider more general spaces than $L^{2}$. On instance is the work [6], which requires the use of the $L^{p}$ space for $p\neq 2$. Another instance is the recent work [23], which requires the use of weighted spaces like the one considered in this paper.
@@
$\|f\|_{2,a}:=\Big{(}\int_{\mathbb{R}^{2[\![n]\!]}}\mathrm{d}x\,\big{|}f(x)e^{a|x|_{1}}\big{|}^{2}\Big{)}^{1/2}\ ,\qquad L^{2}_{a}(\mathbb{R}^{2[\![n]\!]}):=\big{\{}f\,\big{|}\,\|f\|_{2,a}<\infty\big{\}}\ ,$ (1.3)
let $\langle f_{1},f_{2}\rangle:=\int_{\mathbb{R}^{2[\![n]\!]}}\mathrm{d}x\,\overline{f_{1}(x)}f_{2}(x)$ denote the inner product, and for a linear operator $\mathcal{T}:L^{2}_{a}(\mathbb{R}^{2[\![n]\!]})\to L^{2}_{a}(\mathbb{R}^{2[\![n]\!]})$, let
$\|\mathcal{T}\|_{2,a\to 2,a}:=\sup\big{\{}\langle f^{\prime},\mathcal{T}f\rangle\,\big{|}\,\|f^{\prime}\|_{2,a}\leq 1\,,\|f\|_{2,a}\leq 1\big{\}}$ (1.4)
@@
We note in the passing that the work [9] studied a similar problem for bounded $f$s, obtained the pointwise estimates of the prelimiting semigroup (for a different model than (1.1)), and proved the pointwise convergence of that semigroup.
Here is an overview of the proof. As will be seen in Section 2, proving Theorem 1.1 boils down to bounding the weighted norms of the various operators that make up $\mathcal{Q}(t)$ and $\mathcal{Q}^{\varepsilon}(t)$. We obtain these bounds in two steps, by first employing a comparison argument to reduce proving these bounds to proving their unweighted analogs, and then proving their unweighted analogs. While most of the unweighted analogs follow from existing results, one of them, (3.10d), does not seem straightforward and requires significant work. We note that the bounds (3.10a)–(3.10d) are used in the recent work [23] to study the SHF. One may also consider the weighted $L^{p}$ space for $p\neq 2$, and we leave that to future work.
Let us mention a few more works on the moments of the noise-mollified SHE and point to the references in [6] for more references. The work [17] carried out detailed analysis of the second moment and further showed that the pointwise limits of the third and second moments behave inconsistently with the lognormal distribution. The works [7, 11, 12] studied properties related to the Gaussian multiplicative chaos. High moments of the polymers and the mollified SHE are studied in [13, 21, 14].
@@
Acknowledgment
SS is partially supported by the NSF through DMS-2243112. LCT is partially supported by the NSF through DMS-2243112 and the Alfred P. Sloan Foundation through the Sloan Research Fellowship FG-2022-19308.
2. The limiting semigroup, Proof of Theorem 1.1(1)
@@ Acknowledgment
$\mathbb{R}^{2[\![n]\!]}:=(\mathbb{R}^{2})^{[\![n]\!]}$ :=$:=\big{\{}(x_{i})_{i\in[\![n]\!]}\,\big{|}\,x_{i}\in\mathbb{R}^{2}\big{\}}\ ,$ (2.1)
$\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}:=\mathbb{R}^{2}\times\mathbb{R}^{2([\![n]\!]\setminus\alpha)}$ :=$:=\big{\{}y=(y_{\mathrm{c}},(y_{i})_{i\in[\![n]\!]\setminus\alpha})\,\big{|}\,y_{c},y_{i}\in\mathbb{R}^{2}\big{\}}\ ,$ (2.2)
where we index the first coordinate in (2.2) by “c” for “center of mass”. Consider
$S_{\alpha}:\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}\to\mathbb{R}^{2[\![n]\!]},\qquad\big{(}S_{\alpha}y\big{)}_{i}:=\begin{cases}y_{\mathrm{c}}&\text{when }i\in\alpha\\ y_{i}&\text{when }i\in[\![n]\!]\setminus\alpha\ .\end{cases}$ (2.3)
@@
$\mathsf{j}(t):=\int_{0}^{\infty}\mathrm{d}u\frac{t^{u-1}e^{\theta u}}{\Gamma(u)}\ ,$ (2.4)
where $\theta\in\mathbb{R}$ is the parameter in (1.2). For $\alpha\neq\alpha^{\prime}\in\mathrm{Pair}[\![n]\!]$, define the integral operators $\mathcal{P}_{\alpha}(t)$, $\mathcal{P}_{\alpha}(t)^{*}$, $\mathcal{P}_{\alpha\alpha^{\prime}}(t)$, $\mathcal{J}_{\alpha}(t)$ through their kernels as
$\mathcal{P}_{\alpha}(t,y,x)$ $:=\mathcal{P}\big{(}t,S_{\alpha}y-x\big{)}=:\big{(}\mathcal{P}_{\alpha}\big{)}^{*}(t,x,y)\ ,$ (2.5a)
$\mathcal{P}_{\alpha\alpha^{\prime}}(t,y,y^{\prime})$ $:=\mathcal{P}\big{(}t,S_{\alpha}y-S_{\alpha^{\prime}}y^{\prime}\big{)}\ ,$ (2.5b)
$\mathcal{J}_{\alpha}(t,y,y^{\prime})$ $:=4\pi\,\mathsf{j}(t)\,\mathsf{p}(\tfrac{t}{2},y_{\mathrm{c}}-y_{\mathrm{c}}^{\prime})\cdot\prod_{i\in[\![n]\!]\setminus\alpha}\mathsf{p}(t,y_{i}-y^{\prime}_{i})\ ,$ (2.5c)
where $x\in\mathbb{R}^{2[\![n]\!]}$ and $y,y^{\prime}\in\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}$ in (2.5a) and (2.5c), and $y\in\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}$ and $y^{\prime}\in\mathbb{R}^{2[\![n]\!]\setminus\alpha^{\prime}}$ in (2.5b). Next, let
$\mathrm{Dgm}[\![n]\!]:=\big{\{}\vec{\alpha}=(\alpha_{k})_{k=1}^{m}\in\mathrm{Pair}[\![n]\!]^{m}\,\big{|}\,m\in\mathbb{Z}_{>0},\alpha_{k}\neq\alpha_{k+1}\text{ for }k=1,\ldots,m-1\big{\}}\ .$ (2.6)
@@
$\mathcal{R}_{\vec{\alpha}}(t)$ $:=\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\ \mathcal{P}_{\alpha_{1}}(\tau_{\frac{1}{2}})^{*}\prod_{k=1}^{|\vec{\alpha}|-1}\mathcal{J}_{\alpha_{k}}(\tau_{k})\,\mathcal{P}_{\alpha_{k}\alpha_{k+1}}(\tau_{k+\frac{1}{2}})\cdot\mathcal{J}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|})\,\mathcal{P}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|+\frac{1}{2}})\ .$ (2.7)
@@
$\mathcal{Q}(t):=\mathcal{P}(t)+\sum_{\vec{\alpha}\in\mathrm{Dgm}[\![n]\!]}\mathcal{R}_{\vec{\alpha}}(t)\ .$ (2.8)
In [20, Section 8], the series (2.8) is shown to converge in the $L^{2}$ operator norm. Here we seek to do similarly for the weighted operator norm. Let us prepare some notation and tools. First,
$\|f\|_{2,a}^{2}$ $=\|f\|_{L^{2}_{a}(\Omega)}^{2}:=\begin{cases}\int_{\Omega}\mathrm{d}x\,\big{|}f(x)e^{a|x|_{1}}\big{|}^{2}&\Omega=\mathbb{R}^{2[\![n]\!]}\ ,\\ \int_{\Omega}\mathrm{d}y\,\big{|}f(y)e^{2a|y_{\mathrm{c}}|_{1}+a\sum_{i\in[\![n]\!]\setminus\alpha}|y_{i}|_{1}}\big{|}^{2}&\Omega=\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}\ .\end{cases}$ (2.9)
For $\Omega=\mathbb{R}^{2[\![n]\!]}$, the norm is the same as (1.3), but for $\Omega=\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}$, we assign twice the exponential weight to $y_{\mathrm{c}}$, which is natural because it bears a meaning of “merging” the two coordinates in $\alpha$. Accordingly, for operators that map between functions on the spaces (2.1)–(2.2), define the weighted operator norm $\|\cdot\|_{2,a\to 2,a}$ as (1.4) with the $\|\cdot\|_{2,a}$ given in (2.9). We will omit the underlying spaces $\Omega$ when writing the operator norms, because the spaces can be read off from the definition of the operators. For example, referring to (2.5a) and the description that follows, we see that $\|\mathcal{P}_{\alpha}(t)\|_{2,a\to 2,a}$ means the operator norm of $\mathcal{P}_{\alpha}(t):L^{2}_{a}(\mathbb{R}^{2[\![n]\!]})\to L^{2}_{a}(\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha})$. Next, note that the kernels in (2.5) are nonnegative. For a one-parameter family of integral operators $\mathcal{T}(t)$, $t>0$, with a nonnegative kernel $\mathcal{T}(t,z,z^{\prime})$, let
$\Big{\|}\int_{0}^{b}\mathrm{d}t\,\mathcal{T}(t)\Big{\|}_{2,a\to 2,a}:=\sup\Big{\{}\int_{0}^{b}\mathrm{d}t\int\mathrm{d}z^{\prime}\mathrm{d}z\ |f^{\prime}(z^{\prime})|\cdot\mathcal{T}(t,z^{\prime},z)\cdot|f(z)|\Big{\}}\ ,$ (2.10)
where the supremum runs over $\|f^{\prime}\|_{2,a}\leq 1$ and $\|f\|_{2,a}\leq 1$. If this quantity is finite, the operator $\int_{0}^{b}\mathrm{d}t\,\mathcal{T}(t)$ is well-defined and bounded. When $a=0$, the weighted norms reduce to the $L^{2}$ norms, which we write as $\|\cdot\|_{2,0}=\|\cdot\|_{2}$ and $\|\cdot\|_{2,0\to 2,0}=\|\cdot\|_{2\to 2}$. The following lemma will come in handy.
@@
$\big{\|}\mathcal{T}_{\kappa}(t)\big{\|}_{\mathrm{op}}$ $\leq c_{0}e^{c_{0}t}\cdot\begin{cases}t^{-1/2}&\text{when }\kappa=\tfrac{1}{2}\ ,\\ t^{-1}\,\big{|}\log(\tfrac{1}{2}\wedge t)\big{|}^{-2}&\text{when }\kappa\in\mathbb{Z}\cap[1,m]\ ,\\ t^{-1}&\text{when }\kappa\in(\tfrac{1}{2}+\mathbb{Z})\cap(1,m)\ ,\\ t^{-1/2}&\text{when }\kappa=m+\tfrac{1}{2}\ ,\end{cases}$ (2.11)
$\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-c_{0}t}\mathcal{T}_{\kappa}(t)\Big{\|}_{\mathrm{op}}$ $\leq c_{0}\ \text{ when }\kappa\in(\tfrac{1}{2}+\mathbb{Z})\cap(1,m)\ .$ (2.12)

### v2

@@
[…]
1 Introduction
In this paper, we study the two-dimensional, $n$-particle delta-Bose gas. Fix a nonnegative $\Phi\in C^{\infty}(\mathbb{R}^{2})$ with a compact support such that $\int_{\mathbb{R}^{2}}\mathrm{d}x\,\Phi(x)=1$, and, for $\varepsilon\to 0$, let $\Phi^{\varepsilon}(x):=\varepsilon^{-2}\Phi(\varepsilon^{-1}x)$ be the corresponding approximation of the delta function. Let $[\![n]\!]:=\{1,\ldots,n\}$, write $\alpha=ij=\{i<j\}$ for an unordered pair of integers in $[\![n]\!]$, view $ij$ as a set, and let $\mathrm{Pair}[\![n]\!]:=\{ij|ij\subset[\![n]\!]\}$. Consider $\mathbb{R}^{2[\![n]\!]}:=(\mathbb{R}^{2})^{[\![n]\!]}=\{x=(x_{1},\ldots,x_{n})|x_{i}\in\mathbb{R}^{2}\}$, and, for $\alpha=ij$, let $\Phi^{\varepsilon}_{\alpha}(x):=\Phi^{\varepsilon}(x_{i}-x_{j})$. The Hamiltonian $\mathcal{H}^{\varepsilon}$ of the mollified delta-Bose gas acts on functions on $\mathbb{R}^{2[\![n]\!]}$ and is given by
$-\mathcal{H}^{\varepsilon}:=\frac{1}{2}\sum_{i=1}^{n}\Delta_{i}+\beta_{\varepsilon}\sum_{\alpha\in\mathrm{Pair}[\![n]\!]}\Phi^{\varepsilon}_{\alpha}\ ,$ (1)
@@ 1 Introduction
$\beta_{\varepsilon}$ $:=\frac{2\pi}{|\log\varepsilon|}+\frac{\pi}{|\log\varepsilon|^{2}}\Big{(}\theta-2\log 2+2\gamma+2\int_{\mathbb{R}^{4}}\mathrm{d}x\,\mathrm{d}x^{\prime}\Phi(x)\log|x-x^{\prime}|\Phi(x^{\prime})\Big{)}\ ,$ (2)
@@
$\gamma=0.577\ldots$ denotes the Euler–Mascheroni constant, and $\theta$ is a fixed parameter that can be taken to be any real value. In (2), the constants after $\theta$ are just conventional (and can be absorbed into $\theta$).
@@
The delta-Bose gas is an instance of quantum many-body systems and is highly relevant in the study of the Stochastic Heat Equation (SHE). In the quantum context, (1) can be used to approximate the Hamiltonian of $n$ Bosons with pairwise, attractive (true) delta potentials. Quantum systems with delta potentials exhibit intriguing physical properties, pose significant mathematical challenges, and enjoy a long history of study. We point to [2] for a review and mention the works [15, 22, 16, 19, 18] on the two-dimensional delta-Bose gas. As it turns out, in two dimensions, in order to see a non-degenerate limit as $\varepsilon\to 0$, one needs to attenuate the coupling constant $\beta_{\varepsilon}$ logarithmically around the value $2\pi$ as in (2). This was observed in the $n=2$ case in [2, Chapter I.5] and [3]. In the SHE context, the semigroup $\mathcal{Q}^{\varepsilon}(t):=e^{-t\mathcal{H}^{\varepsilon}}$ describes the moments of the noise-mollified SHE. Based on [2, 1], the work [3] established the convergence (as $\varepsilon\to 0$) of the second moment of the noise-mollified SHE to an explicit limit. Convergence of higher moments are significantly more challenging, and was later obtained in [5] for the third moment based on [4], and in [20] for of all moments based on [22, 16]; see [8, 9, 10] for probabilistic studies of the moments. A central question in this context concerns the $\varepsilon\to 0$ limit of the noise-mollified SHE, and the limit is called the Stochastic Heat Flow (SHF). The construction of the SHF had been a challenging open problem and was accomplished [6].
@@
In this paper, we generalize the results of [20] from the $L^{2}$ space to a weighted $L^{2}$ space. Based on [22, 16], the work [20] gave an explicit formula of a strongly continuous semigroup $\mathcal{Q}(t)$ on $L^{2}(\mathbb{R}^{2[\![n]\!]})$ and showed that $\mathcal{Q}^{\varepsilon}(t):=e^{-t\mathcal{H}^{\varepsilon}}$ converges strongly to $\mathcal{Q}(t)$ on $L^{2}(\mathbb{R}^{2[\![n]\!]})$. These results translate into the convergence of the moments of the noise-mollified SHE to an explicit limit, for $L^{2}$ initial conditions. The class of $L^{2}$ functions is natural, because the generator of $\mathcal{Q}(t)$ is self-adjoint, but the class leaves out some commonly considered initial conditions, most notably the constant/flat ones. Further, when using delta-Bose gas to study the noise-mollified SHE and the SHF, one often needs to consider more general spaces than $L^{2}$. One instance is the work [6], which introduces weighted $L^{p}$ estimates for the semigroup in a discrete setting. Another instance is the recent work [23], which requires the use of the weighted $L^{2}$ space and the bounds (29)–(30) from this paper.
@@
$\|f\|_{2,a}:=\Big{(}\int_{\mathbb{R}^{2[\![n]\!]}}\mathrm{d}x\,\big{|}f(x)e^{a|x|_{1}}\big{|}^{2}\Big{)}^{1/2}\ ,\qquad L^{2}_{a}(\mathbb{R}^{2[\![n]\!]}):=\big{\{}f\,\big{|}\,\|f\|_{2,a}<\infty\big{\}}\ ,$ (3)
let $\langle f_{1},f_{2}\rangle:=\int_{\mathbb{R}^{2[\![n]\!]}}\mathrm{d}x\,e^{2a|x|_{1}}\overline{f_{1}(x)}f_{2}(x)$ denote the inner product, and for a linear operator $\mathcal{T}:L^{2}_{a}(\mathbb{R}^{2[\![n]\!]})\to L^{2}_{a}(\mathbb{R}^{2[\![n]\!]})$, let
$\|\mathcal{T}\|_{2,a\to 2,a}:=\sup\big{\{}\langle f^{\prime},\mathcal{T}f\rangle\,\big{|}\,\|f^{\prime}\|_{2,a}\leq 1\,,\|f\|_{2,a}\leq 1\big{\}}$ (4)
@@
We note in the passing that the work [9] studied a similar problem for bounded $f$s and proved the pointwise (in $x$) convergence of $(\mathcal{Q}^{\varepsilon}(t)f)(x)$.
A crucial property used in [20] is that the generators of $\mathcal{Q}(t)$ and $\mathcal{Q}^{\varepsilon}(t)$ are self-adjoint in $L^{2}$, but this property no longer holds in $L^{2}_{a}$. In the self-adjoint setting of [20], to prove statements like those in Theorem 1.1, one can just work with the resolvents. Indeed, in the self-adjoint setting, thanks to the spectral theorem, bounds on the resolvents imply bounds on the spectra and hence the semigroups. In our non-self-adjoint setting, we need to bound the semigroups $\mathcal{Q}(t)$ and $\mathcal{Q}^{\varepsilon}(t)$ directly, without resorting to resolvents. These semigroups are much less well-behaved than their resolvents, so the bounds require additional work. This fact manifests itself in our proof of (30d).
Here is an overview of the proof. As will be seen in Section 2, proving Theorem 1.1 boils down to bounding the weighted norms of the various operators that make up $\mathcal{Q}(t)$ and $\mathcal{Q}^{\varepsilon}(t)$. We obtain these bounds in two steps, by first employing a comparison argument to reduce proving these bounds to proving their unweighted analogs, and then proving their unweighted analogs. While most of the unweighted analogs follow from existing results, one of them, (30d), does not seem straightforward and requires significant work. One may also consider the weighted $L^{p}$ space for $p\neq 2$, and we leave that to future work.
Let us mention a few more works on the moments of the noise-mollified SHE and point to the references in [6] for more references. The work [17] carried out detailed analysis of the second moment and further used of the Gaussian correlation inequality to compare the moments of the SHE with those of a log-normal. The works [7, 11, 12] studied properties related to the Gaussian multiplicative chaos. High moments of the polymers and the mollified SHE are studied in [13, 21, 14].
@@
2 The limiting semigroup, Proof of Theorem 1.1(1)
@@
$\mathbb{R}^{2[\![n]\!]}:=(\mathbb{R}^{2})^{[\![n]\!]}$ :=$:=\big{\{}(x_{i})_{i\in[\![n]\!]}\,\big{|}\,x_{i}\in\mathbb{R}^{2}\big{\}}\ ,$ (5)
$\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}:=\mathbb{R}^{2}\times\mathbb{R}^{2([\![n]\!]\setminus\alpha)}$ :=$:=\big{\{}y=(y_{\mathrm{c}},(y_{i})_{i\in[\![n]\!]\setminus\alpha})\,\big{|}\,y_{c},y_{i}\in\mathbb{R}^{2}\big{\}}\ ,$ (6)
where we index the first coordinate in (6) by “c” for “center of mass”. Consider
$S_{\alpha}:\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}\to\mathbb{R}^{2[\![n]\!]},\qquad\big{(}S_{\alpha}y\big{)}_{i}:=\begin{cases}y_{\mathrm{c}}&\text{when }i\in\alpha\\ y_{i}&\text{when }i\in[\![n]\!]\setminus\alpha\ .\end{cases}$ (7)
@@
$\mathsf{j}(t):=\int_{0}^{\infty}\mathrm{d}u\frac{t^{u-1}e^{\theta u}}{\Gamma(u)}\ ,$ (8)
where $\theta\in\mathbb{R}$ is the parameter in (2). For $\alpha\neq\alpha^{\prime}\in\mathrm{Pair}[\![n]\!]$, define the integral operators $\mathcal{P}_{\alpha}(t)$, $\mathcal{P}_{\alpha}(t)^{*}$, $\mathcal{P}_{\alpha\alpha^{\prime}}(t)$, $\mathcal{J}_{\alpha}(t)$ through their kernels as
$\mathcal{P}_{\alpha}(t,y,x)$ $:=\mathcal{P}\big{(}t,S_{\alpha}y-x\big{)}=:\big{(}\mathcal{P}_{\alpha}\big{)}^{*}(t,x,y)\ ,$ (9a)
$\mathcal{P}_{\alpha\alpha^{\prime}}(t,y,y^{\prime})$ $:=\mathcal{P}\big{(}t,S_{\alpha}y-S_{\alpha^{\prime}}y^{\prime}\big{)}\ ,$ (9b)
$\mathcal{J}_{\alpha}(t,y,y^{\prime})$ $:=4\pi\,\mathsf{j}(t)\,\mathsf{p}(\tfrac{t}{2},y_{\mathrm{c}}-y_{\mathrm{c}}^{\prime})\cdot\prod_{i\in[\![n]\!]\setminus\alpha}\mathsf{p}(t,y_{i}-y^{\prime}_{i})\ ,$ (9c)
where $x\in\mathbb{R}^{2[\![n]\!]}$ and $y,y^{\prime}\in\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}$ in (9a) and (9c), and $y\in\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}$ and $y^{\prime}\in\mathbb{R}^{2[\![n]\!]\setminus\alpha^{\prime}}$ in (9b). Next, let
$\mathrm{Dgm}[\![n]\!]:=\big{\{}\vec{\alpha}=(\alpha_{k})_{k=1}^{m}\in\mathrm{Pair}[\![n]\!]^{m}\,\big{|}\,m\in\mathbb{Z}_{>0},\alpha_{k}\neq\alpha_{k+1}\text{ for }k=1,\ldots,m-1\big{\}}\ .$ (10)
@@
$\mathcal{R}_{\vec{\alpha}}(t)$ $:=\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\ \mathcal{P}_{\alpha_{1}}(\tau_{\frac{1}{2}})^{*}\prod_{k=1}^{|\vec{\alpha}|-1}\mathcal{J}_{\alpha_{k}}(\tau_{k})\,\mathcal{P}_{\alpha_{k}\alpha_{k+1}}(\tau_{k+\frac{1}{2}})\cdot\mathcal{J}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|})\,\mathcal{P}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|+\frac{1}{2}})\ .$ (11)
@@
$\mathcal{Q}(t):=\mathcal{P}(t)+\sum_{\vec{\alpha}\in\mathrm{Dgm}[\![n]\!]}\mathcal{R}_{\vec{\alpha}}(t)\ .$ (12)
In [20, Section 8], the series (12) is shown to converge in the $L^{2}$ operator norm. Here we seek to do similarly for the weighted operator norm. Let us prepare some notation and tools. First,
$\|f\|_{2,a}^{2}$ $=\|f\|_{L^{2}_{a}(\Omega)}^{2}:=\begin{cases}\int_{\Omega}\mathrm{d}x\,\big{|}f(x)e^{a|x|_{1}}\big{|}^{2}&\Omega=\mathbb{R}^{2[\![n]\!]}\ ,\\ \int_{\Omega}\mathrm{d}y\,\big{|}f(y)e^{2a|y_{\mathrm{c}}|_{1}+a\sum_{i\in[\![n]\!]\setminus\alpha}|y_{i}|_{1}}\big{|}^{2}&\Omega=\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}\ .\end{cases}$ (13)
For $\Omega=\mathbb{R}^{2[\![n]\!]}$, the norm is the same as (3), but for $\Omega=\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}$, we assign twice the exponential weight to $y_{\mathrm{c}}$, which is natural because it bears a meaning of “merging” the two coordinates in $\alpha$. Accordingly, for operators that map between functions on the spaces (5)–(6), define the weighted operator norm $\|\cdot\|_{2,a\to 2,a}$ as (4) with the $\|\cdot\|_{2,a}$ given in (13). We will omit the underlying spaces $\Omega$ when writing the operator norms, because the spaces can be read off from the definition of the operators. For example, referring to (9a) and the description that follows, we see that $\|\mathcal{P}_{\alpha}(t)\|_{2,a\to 2,a}$ means the operator norm of $\mathcal{P}_{\alpha}(t):L^{2}_{a}(\mathbb{R}^{2[\![n]\!]})\to L^{2}_{a}(\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha})$. Next, note that the kernels in (9) are nonnegative. For a one-parameter family of integral operators $\mathcal{T}(t)$, $t>0$, with a nonnegative kernel $\mathcal{T}(t,z,z^{\prime})$, if
$\sup\Big{\{}\int_{0}^{b}\mathrm{d}t\int\mathrm{d}z^{\prime}\mathrm{d}z\ |f^{\prime}(z^{\prime})|\cdot\mathcal{T}(t,z^{\prime},z)\cdot|f(z)|\Big{\}}\ ,$ (14)
is finite when the supremum is taken over $\|f^{\prime}\|_{2,a}\leq 1$ and $\|f\|_{2,a}\leq 1$, then the operator $\int_{0}^{b}\mathrm{d}t\,\mathcal{T}(t)$ is well-defined and bounded with $\big{\|}\int_{0}^{b}\mathrm{d}t\,\mathcal{T}(t)\big{\|}_{2,a\to 2,a}$ equal to the above expression. When $a=0$, the weighted norms reduce to the $L^{2}$ norms, which we write as $\|\cdot\|_{2,0}=\|\cdot\|_{2}$ and $\|\cdot\|_{2,0\to 2,0}=\|\cdot\|_{2\to 2}$. It will be useful to be able to bound convolution-like integral as in equation (11) in norm by bounding the individual integrands. To this end, we have the following lemma:
@@
$\big{\|}\mathcal{T}_{\kappa}(t)\big{\|}_{\mathrm{op}}$ $\leq c_{0}e^{c_{0}t}\cdot\begin{cases}t^{-1/2}&\text{when }\kappa=\tfrac{1}{2}\ ,\\ t^{-1}\,\big{|}\log(\tfrac{1}{2}\wedge t)\big{|}^{-2}&\text{when }\kappa\in\mathbb{Z}\cap[1,m]\ ,\\ t^{-1}&\text{when }\kappa\in(\tfrac{1}{2}+\mathbb{Z})\cap(1,m)\ ,\\ t^{-1/2}&\text{when }\kappa=m+\tfrac{1}{2}\ ,\end{cases}$ (15)
$\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-c_{0}t}\mathcal{T}_{\kappa}(t)\Big{\|}_{\mathrm{op}}$ $\leq c_{0}\ \text{ when }\kappa\in(\tfrac{1}{2}+\mathbb{Z})\cap(1,m)\ .$ (16)